The larger of two numbers is six more than the smaller. If the larger is doubled and the smaller is increased by 24. find both the numbers
Answer by mananth(16949)

The larger of two numbers is six more than the smaller. If the larger is doubled and the smaller is increased by 24. find both the numbers
..
smaller = x
larger = x+6
..
2(x+6)=x+24
2x+12=x+24
2x-x=24-12
x=12
..
12 & 18 are the numbers
..
m.ananth@hotmail.ca